OpenStudy (anonymous):

A package contains 18 candy canes, 13 of which are cracked. If 4 are selected, find the probability of getting exactly four cracked candy canes.

hero (hero):

Hint: The probability of selecting one cracked candy cane is \(\dfrac{13}{18}\) The probability of selecting a second cracked candy cane is \(\dfrac{13 - 1}{18 - 1}\) The probability of selecting a third cracked candy cane is \(\dfrac{13 - 2}{18 - 2}\) Notice a pattern? After finding the probability of selecting the fourth cracked candy cane, multiply all four probabilities together to find the probability of selecting four cracked candy canes